# A Two-Timescale Primal-Dual Framework for Reinforcement Learning via Online Dual Variable Guidance

### v2

@@ 1 Introduction
Recent years have seen many new promising developments in the theory and application of reinforcement learning (RL) as a framework for sequential decision making with notable successes of regularization-based algorithms, for example, the Soft Actor-Critic [Haarnoja et al., 2018] or Trust Region Policy Optimization [Schulman et al., 2015]. The regularization typically takes the form of a convex penalty term in the policy, which is a setting that has been subsequently formalized by Geist et al. [2019], Neu et al. [2017]. The adjustment of the objective aims at inducing desirable exploratory behavior of the learned policies and stability in their training [Schulman et al., 2017], and allows the application of dynamical systems analysis that takes advantage of the adjusted problem’s strongly convex structure [Li et al., 2024].
@@
As another development, the linear programming (LP) reformulation of Markov Decision Processes (MDPs), originally dating back to Manne [1960], Hernández-Lerma and Lasserre [1996], Borkar [2002], has recently received renewed interest. It has led to LP-based algorithms shown to be on par with broadly applied benchmarks, for example, the Dual Actor-Critic [Dai et al., 2018a]. The LP approach to RL allows the reformulation of the Bellman optimality equation, which characterizes an optimal policy, as a min-max saddle-point problem that can be addressed with gradient methods. The analysis of the various algorithms proposed in this line of research often relies on simplifying assumptions such as access to a sampling generator [Chen and Wang, 2016], a known model [Li et al., 2024], or the uniqueness of the optimal policy [Lee and He, 2019]. A further challenge in solving the min-max RL problem formulation lies in the nested loop structure that arises when the inner optimization is first solved approximately during each outer iteration, which complicates the convergence analysis [Gabbianelli et al., 2024, Dai et al., 2018a].
@@
The asymptotic convergence under general model-access assumptions is well understood for (approximate) dynamic programming algorithms; (see Tsitsiklis [1994], Jaakkola et al. [1993]). However, finite-time guarantees for the last iterate remain scarce for LP-based primal-dual schemes driven by Markovian single-trajectory data and iterate-dependent (on-policy) exploration, particularly when gradients are formed from replay buffers and are therefore biased at finite times. This perspective is complementary to regret-minimization in online RL: rather than optimizing regret or minimax sample complexity via exploration bonuses, we study the stability and last-iterate behavior of a primalreplay-buffer biasdual stochastic approximation method for an LP saddle-point formulation in a realistically correlated-data regime.
@@
In this work, we combine recent advances on regularized LP formulations of RL [Li et al., 2024] with the classical theory of stochastic approximation [Borkar, 2023] and modern finite-time analyses for two-timescale stochastic approximation [Zeng et al., 2024]. Our goal is to design a provably convergent algorithm that maintains on-policy exploration while performing off-policy updates using previously observed transitions stored in an experience replay buffer, and to quantify its last-iterate finite-time rate. The main contributions of this paper are summarized as follows:
@@
Dai et al. [2018a] introduced a dual actor-critic algorithm rooted in the LP formulation that incorporates multi-step Bellman equations, path regularization, and a nested stochastic dual-ascent update. While their approach shares similarities with ours in starting from an LP/Lagrangian viewpoint and employing convexity-inducing regularizers, we adopt a two-timescale stochastic approximation framework that yields an incremental update rule and enables an almost sure convergence analysis under Markovian sampling. Similarly, works by Chen and Wang [2016], Wang [2017], Chen et al. [2018] study LP formulations for discounted, average reward, and finite-horizon MDPs through projected stochastic primal-dual methods (including incremental coordinate updates). These works provide finite-sample or asymptotic convergence results under the crucial assumption of access to a sampling generator. In contrast, we focus on the last iterate and establish almost sure convergence assuming only access to a single trajectory of state-action pairs (Markovian model access). Lee and He [2019] extends Chen and Wang [2016] by incorporating a Q-function estimation step. Their analysis allows time-varying behavioral policies under conditions requiring sublinear convergence of the induced state-action distribution, which can be difficult to verify a priori and makes fully on-policy analysis nontrivial. Moreover, their algorithm relies on samples drawn from the stationary state-action distribution of the current policy. In Bas-Serrano et al. [2021] and the subsequent thesis Bas Serrano [2022], the authors introduce the logistic Bellman error derived from an entropy-regularized LP formulation and establish finite-sample rates in the tabular setting under stationary-distribution generator access. In comparison, our stochastic approximation analysis further allows us to avoid stationary distribution sampling and to analyze iterate-dependent (on-policy) exploration under Markovian single-trajectory access. Under a strengthened ergodicity condition on the Markov chain induced by the iterate-dependent behavioral policy, we additionally establish a last-iterate finite-time mean-square convergence rate.
Two-timescale stochastic approximation has been applied to dynamic programming-based actor-critic algorithms [Konda and Borkar, 1999] and to constrained MDPs [Borkar, 2005], where safety constraints are added to the reward maximization objective. The Lagrangian relaxation of the constrained optimization problem motivates the application of various primal-dual methods [Chen et al., 2024, Hong and Tewari, 2024, Li et al., 2023].
The LP approach has also been applied to Offline RL problems, where near-optimal policies are learned from static datasets of transitions [Zhan et al., 2022, Ozdaglar et al., 2023, Gabbianelli et al., 2024]. Even though the analysis objective differs, since we consider the online setting, their saddle-point problem resulting from the Lagrangian dual formulation of the LP is closely related to our optimization target.
Parallel investigations into regularized MDPs have also been influential. Li et al. [2024] explore a primal-dual formulation of entropy-regularized MDPs with additional convexity in the primal variable, providing asymptotic convergence analysis using natural gradient methods, albeit under the assumption of full model knowledge. Neu et al. [2017] and Geist et al. [2019] formalize the entropy-regularized MDP framework, outlining the conditional entropy regularizer and the contraction properties of the regularized Bellman evaluation operator. Ying and Zhu [2020] offers a comprehensive overview of various LP formulations in both regularized and unregularized contexts, emphasizing their equivalence to Bellman equations and policy gradient methods. Furthermore, convex Q-learning [Meyn, 2022, Mehta and Meyn, 2020, Lu and Meyn, 2023] is motivated by the idea of deriving a convex relaxation of the squared Bellman error based on the LP approach. The method involves solving a constrained convex optimization problem over the primal variable, the Q-function. It is primarily designed for analyzing the non-tabular setting, where the value function is learned via function approximation, which distinguishes this approach from our own.
Finally, our finite-time analysis builds on recent last-iterate convergence-rate results for two-timescale stochastic approximation Zeng et al. [2024]. We face additional challenges in our setting, namely asynchronous updates and replay-buffer gradient estimates that introduce bias. Despite this, on the high-probability good visitation event $\mathcal{G}_{\delta}$ (after a burn-in period), we prove a last-iterate mean-square rate $\widetilde{\mathcal{O}}(k^{-2/3})$ for the dual iterates, implying corresponding mean-square rates for the induced regularized value function and policy. Equivalently, achieving $\epsilon$ accuracy in mean-square error requires $k(\epsilon)=\widetilde{\mathcal{O}}(\epsilon^{-3/2})$ iterations to ensure $\mathbb{E}[\|\rho_{k}-\rho^{\star}\|_{2}^{2}\mid\mathcal{G}_{\delta}]\leq\epsilon$ (up to logarithmic factors).
@@
Chen and Wang [2016] Generator Tabular PAC Duality gap $\widetilde{\mathcal{O}}(\epsilon^{-2})$
Dai et al. [2018a] Markovian Non-linear Sketch
Lee and He [2019] Stationary Distr. Generator Tabular PAC Duality gap $\widetilde{\mathcal{O}}(\epsilon^{-2})$
Bas Serrano [2022] Stationary Distr. Generator Tabular PAC $\epsilon$-optimal policy $\widetilde{\mathcal{O}}(\epsilon^{-9})$
Gabbianelli et al. [2024] Offline Linear PAC $\epsilon$-optimal policy $\widetilde{\mathcal{O}}(\epsilon^{-2})$
Li et al. [2024] Full model Tabular Deterministic asymptotic
@@
  • Table 1: Overview of LP-based RL methods and their guarantees. We summarize related approaches based on linear-programming or Lagrangian formulations of MDPs, organized by the model access assumption, the function-approximation regime, the type of analysis, and the corresponding guarantee and reported rate. The “Rate” column is stated with respect to the guarantee listed in the preceding column, and is therefore not directly comparable across rows when the guarantees or access models differ.
Structure.
@@
The structure of the paper is as follows. Section 2 formally introduces regularized MDPs and the associated saddle-point formulation that underlies our proposed algorithms. Section 3 presents a first algorithm in a synchronous setting, assuming access to a generative model. In Section 4, we extend this setting to the asynchronous case, where learning occurs from a single trajectory of data generated under a potentially changing policy. The section further contains the almost sure convergence result and our convergence rate analysis. Section 5 demonstrates the effectiveness of the proposed asynchronous algorithm on a standard reinforcement learning benchmark. Finally, Section 6 concludes the paper. The proofs details and technical results are deferred to the appendix to improve the readability of the main text.
Notation.
@@
For any state $s\in\mathcal{S}$, we denote by $\delta_{s}$ the Dirac measure in $s$. For a variable $v(s,a)$ depending on state-action pairs $(s,a)\in\mathcal{S}\times\mathcal{A}$, we denote the state marginal by $\tilde{v}(s):=\sum_{a\in\mathcal{A}}v(s,a)$. For a finite set $\mathcal{S}$ with cardinality $n$, the probability simplex over $\mathcal{S}$ is denoted by $\Delta_{\mathcal{S}}=\{x\in\mathbb{R}_{+}^{n}:\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}x_{i}=1\}$. We use the symbol $e_{i}\in\mathbb{R}^{|\mathcal{S}|}$ for the $i$-th standard basis vector on $\mathbb{R}^{|\mathcal{S}|}$. For any logical expression $\mathcal{E}$, the indicator function $\mathbf{1}\{\mathcal{E}\}$ evaluates to $1$ if $\mathcal{E}$ is true and to 0 otherwise. To denote sequences, we use $\{\beta_{k}\}$ and $\{\beta(k)\}$ interchangeably for $k\in\mathbb{N}$. For $a,b\in\mathbb{R}$ we denote with $a\lor b$ the larger of $a$ and $b$ and with $\lfloor a\rfloor$ the integer part of $a$. For a compact set $H\subset\mathbb{R}^{d}$, $d\in\mathbb{N}$, we denote the Euclidean projection operator as $\Pi_{H}$, where $\Pi_{H}:\mathbb{R}^{d}\rightarrow H$ is defined by $\Pi_{H}(x)=\arg\min_{y\in H}\|x-y\|_{2}$. We denote the total variation norm as $\|\cdot\|_{\mathrm{TV}}$. For kernels $\mathcal{P},\,\mathcal{Q}$ on $\mathcal{X}$, we define $\|\mathcal{P}-\mathcal{Q}\|_{\infty}\;:=\;\sup_{x\in\mathcal{X}}\,\|\mathcal{P}(\cdot|x)-\mathcal{Q}(\cdot|x)\|_{\mathrm{TV}}.$
@@ 2 Problem Formulation
which admits a unique solution. We approach the unregularized MDP by its formulation as a linear program, Puterman [1994, Section 6.9]. Reformulating the Bellman optimality equations as inequalities yields that a bounded function $V:\mathcal{S}\rightarrow\mathbb{R}$ that satisfies
@@
where $\mu\in\mathbb{R}_{+}^{|\mathcal{S}|}$ is any strictly positive vector such that $\sum_{s\in\mathcal{S}}\mu(s)=1$. It is well-known [Puterman, 1994, Section 6.9] that the optimizer to the primal LP (1) is exactly the optimal value function $V^{\star}_{ur}$. The corresponding dual linear problem is defined as
@@
where $\tilde{\rho}(s):=\sum_{a^{\prime}\in\mathcal{A}}\rho(s,a^{\prime})$ denotes the state marginal of the dual variable. Let $\rho^{\star}_{ur}$ denote a solution to the dual problem (2), then $\pi_{\rho_{ur}^{\star}}$ is an optimal policy, that is, $V^{\pi_{\rho^{\star}_{ur}}}_{ur}=V^{\star}_{ur}$, see [Puterman, 1994, Theorem 6.9.4]. The Lagrangian formulation of the primal LP (1) is
@@ 2.1 Regularized MDP
Our approach is based on the regularized MDP formulation as formally described in Geist et al. [2019]. Let $G:\mathbb{R}^{|\mathcal{A}|}_{+}\rightarrow\mathbb{R}$ be a strongly concave function and $\eta_{\rho}>0$ a regularization parameter. For a stationary policy $\pi\in\Pi$, the regularized Bellman operator $T_{\pi,G}:\mathbb{R}^{|\mathcal{S}|}\rightarrow\mathbb{R}^{|\mathcal{S}|},V\mapsto T_{\pi,G}V$ is defined state-wise as
@@
where the maximum is state-wise. Both operators are $\gamma$-contractions in the supremum norm, [Geist et al., 2019, Prop. 2]. The regularized value function of a policy $\pi$, denoted $V_{r}^{\pi}$, is defined as the fixed point of the operator $T_{\pi,G}$. The regularized value function of the regularized MDP, denoted $V^{\star}_{r}$, is defined as the fixed point of the regularized optimality operator $T_{\star,G}$. Due to the strong concavity of $G$, the regularized Bellman optimality operator has a unique maximizing argument.
Proposition 2.1.
[Geist et al., 2019, Theorem 1] The policy $\pi^{\star}_{r}:=\arg\max_{\pi\in\Pi}T_{\pi,G}V^{\star}_{r}$ is the unique optimal regularized policy in the sense that, state-wise, $V^{\pi^{\star}_{r}}_{r}=V^{\star}_{r}\geq V^{\pi}_{r}$ for all policies $\pi\in\Pi$.
For the Shannon entropy regularizer $G(\pi(\cdot|s)):=\sum_{a\in\mathcal{A}}\pi(a|s)\log(\pi(a|s)),$ we have $0\leq G(\pi(\cdot|s))\leq\log|\mathcal{A}|,$ for all $\pi\in\Pi$ and we can take $L_{G}=0,\ U_{G}=\log|\mathcal{A}|$ in the above stated bounds. The optimal regularized policy $\pi^{\star}_{r}$ is a strictly positive and exploratory softmax policy for entropy-type $G$. The Lagrangian formulation of the regularized MDP [Ying and Zhu, 2020] includes the strongly concave function of the dual variable $g(\rho):=\sum_{s\in\mathcal{S}}G(\rho_{s})$ with $\rho_{s}=\rho(s,\cdot)\in\mathbb{R}^{|\mathcal{A}|}_{+}$ and is given as
@@
The following corollary is a consequence of Proposition 2.1 and characterizes the solution to the regularized Lagrangian.
Corollary 2.1.
@@
Proposition 2.2.
[Geist et al., 2019, Theorem 2] Assume that constants $L_{G}$ and $U_{G}$ exist such that for all $\pi\in\Pi$ we have $L_{G}\leq\max_{s\in\mathcal{S}}G(\pi(\cdot|s))\leq U_{G}$. Then, $V^{\star}_{ur}(s)-\eta_{\rho}\frac{U_{G}-L_{G}}{1-\gamma}\leq V_{ur}^{\pi^{\star}_{r}}(s)\leq V^{\star}_{ur}(s)$ for all $s\in\mathcal{S}$.
To ensure better computational properties, we seek to make the regularized problem (7), defined by the regularized Lagrangian $L_{r}$, strongly convex in $V$. Since $L_{r}$ is originally linear in $V$, we propose an additional modification to problem (7). Specifically, we replace the linear term $\mu^{\top}V$ with a quadratic term and introduce a specific dual regularization function $g$ of the form
@@
The closely related recent work Li et al. [2024] also studies the convexified regularized MDP (9). They prove that the convex modification does not affect the optimal regularized value and only changes the optimal dual variable by a scaling factor that does not influence the optimal regularized policy.
Proposition 2.3.
[Li et al., 2024, Theorem 2.1] The unique solution $(V^{\star},\rho^{\star})$ to the double-regularized problem (9) is such that $V^{\star}=V^{\star}_{r}$ and the optimal policy induced by $\rho^{\star}$ coincides with $\pi^{\star}_{r}$, that is, $\pi_{\rho^{\star}}=\pi_{\rho_{r}^{\star}}=\pi^{\star}_{r}$.
@@
Lemma 2.1 (Optimal dual variable).
The optimal dual variable of the double-regularized problem (9) with bounded reward is strictly positive and bounded from above, that is, there exist $C^{L},C^{U}>0$ such that $C^{L}<\rho^{\star}(s,a)<C^{U}$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$.
Let $C^{L}$ and $C^{U}$ be the explicit bounds derived in the constructive proof of Lemma 2.1. Define $H$ as

### v3

@@ 1 Introduction
Recent years have seen many new promising developments in the theory and application of reinforcement learning (RL) as a framework for sequential decision making with notable successes of regularization-based algorithms, for example, the Soft Actor-Critic (Haarnoja et al., 2018) or Trust Region Policy Optimization (Schulman et al., 2015). The regularization typically takes the form of a convex penalty term in the policy, which is a setting that has been subsequently formalized by Geist et al. (2019); Neu et al. (2017). The adjustment of the objective aims at inducing desirable exploratory behavior of the learned policies and stability in their training (Schulman et al., 2017), and allows the application of dynamical systems analysis that takes advantage of the adjusted problem’s strongly convex structure (Li et al., 2024).
@@
As another development, the linear programming (LP) reformulation of Markov decision processes (MDPs), originally dating back to Manne (1960); Hernández-Lerma and Lasserre (1996); Borkar (2002), has recently received renewed interest. It has led to LP-based algorithms shown to be on par with broadly applied benchmarks, for example, the Dual Actor-Critic (Dai et al., 2018a). The LP approach to RL allows the reformulation of the Bellman optimality equation, which characterizes an optimal policy, as a min-max saddle-point problem that can be addressed with gradient methods. The analysis of the various algorithms proposed in this line of research often relies on simplifying assumptions such as access to a sampling generator (Chen and Wang, 2016), a known model (Li et al., 2024), or the uniqueness of the optimal policy (Lee and He, 2019). A further challenge in solving the min-max RL problem formulation lies in the nested loop structure that arises when the inner optimization is first solved approximately during each outer iteration, which complicates the convergence analysis (Gabbianelli et al., 2024; Dai et al., 2018a).
@@
The asymptotic convergence under general model-access assumptions is well understood for (approximate) dynamic programming algorithms; (see Tsitsiklis (1994); Jaakkola et al. (1993)). However, finite-time guarantees for the last iterate remain scarce for LP-based primal-dual schemes driven by Markovian single-trajectory data and iterate-dependent (on-policy) exploration, particularly when gradients are formed from replay buffers and are therefore biased at finite times. This perspective is complementary to regret-minimization in online RL: rather than optimizing regret or minimax sample complexity via exploration bonuses, we study the stability and last-iterate behavior of a primalreplay-buffer biasdual stochastic approximation method for an LP saddle-point formulation in a realistically correlated-data regime.
@@
In this work, we combine recent advances on regularized LP formulations of RL (Li et al., 2024) with the classical theory of stochastic approximation (Borkar, 2023) and modern finite-time analyses for two-timescale stochastic approximation (Zeng et al., 2024). Our goal is to design a provably convergent algorithm that maintains on-policy exploration while performing off-policy updates using previously observed transitions stored in an experience replay buffer, and to quantify its last-iterate finite-time rate. The main contributions of this paper are summarized as follows:
@@ 1.1 Related Work
Dai et al. (2018a) introduced a dual actor-critic algorithm rooted in the LP formulation that incorporates multi-step Bellman equations, path regularization, and a nested stochastic dual-ascent update. While their approach shares similarities with ours in starting from an LP/Lagrangian viewpoint and employing convexity-inducing regularizers, we adopt a two-timescale stochastic approximation framework that yields an incremental update rule and enables an almost sure convergence analysis under Markovian sampling. Similarly, works by Chen and Wang (2016); Wang (2017); Chen et al. (2018) study LP formulations for discounted, average reward, and finite-horizon MDPs through projected stochastic primal-dual methods (including incremental coordinate updates). These works provide finite-sample or asymptotic convergence results under the crucial assumption of access to a sampling generator. In contrast, we focus on the last iterate and establish almost sure convergence assuming only access to a single trajectory of state-action pairs (Markovian model access). Lee and He (2019) extends Chen and Wang (2016) by incorporating a Q-function estimation step. Their analysis allows time-varying behavioral policies under conditions requiring sublinear convergence of the induced state-action distribution, which can be difficult to verify a priori and makes fully on-policy analysis nontrivial. Moreover, their algorithm relies on samples drawn from the stationary state-action distribution of the current policy. In Bas-Serrano et al. (2021) and the subsequent thesis Bas-Serrano (2022), the authors introduce the logistic Bellman error derived from an entropy-regularized LP formulation and establish finite-sample rates in the tabular setting under stationary-distribution generator access. In comparison, our stochastic approximation analysis further allows us to avoid stationary distribution sampling and to analyze iterate-dependent (on-policy) exploration under Markovian single-trajectory access. Under a strengthened ergodicity condition on the Markov chain induced by the iterate-dependent behavioral policy, we additionally establish a last-iterate finite-time mean-square convergence rate.
@@
Chen and Wang (2016) Generator Tabular PAC Duality gap $\widetilde{\mathcal{O}}(\epsilon^{-2})$
Dai et al. (2018a) Markovian Non-linear Sketch
Lee and He (2019) Stationary Distr. Generator Tabular PAC Duality gap $\widetilde{\mathcal{O}}(\epsilon^{-2})$
Bas-Serrano (2022) Stationary Distr. Generator Tabular PAC $\epsilon$-optimal policy $\widetilde{\mathcal{O}}(\epsilon^{-9})$
Gabbianelli et al. (2024) Offline Linear PAC $\epsilon$-optimal policy $\widetilde{\mathcal{O}}(\epsilon^{-2})$
Li et al. (2024) Full model Tabular Deterministic asymptotic
@@
Table 2: Overview of LP-based RL methods and their guarantees. We summarize related approaches based on linear-programming or Lagrangian formulations of MDPs, organized by the model access assumption, the function-approximation regime, the type of analysis, and the corresponding guarantee and reported rate. The “Rate” column is stated with respect to the corresponding guarantee, and is therefore not directly comparable across rows when the guarantees or access models differ.
Two-timescale stochastic approximation has been applied to dynamic programming-based actor-critic algorithms (Konda and Borkar, 1999) and to constrained MDPs (Borkar, 2005), where safety constraints are added to the reward maximization objective. The Lagrangian relaxation of the constrained optimization problem motivates the application of various primal-dual methods (Chen et al., 2024; Hong and Tewari, 2024; Li et al., 2023).
The LP approach has also been applied to Offline RL problems, where near-optimal policies are learned from static data sets of transitions (Zhan et al., 2022; Ozdaglar et al., 2023; Gabbianelli et al., 2024). Even though the analysis objective differs, since we consider the online setting, their saddle-point problem resulting from the Lagrangian dual formulation of the LP is closely related to our optimization target.
Parallel investigations into regularized MDPs have also been influential. Li et al. (2024) explore a primal-dual formulation of entropy-regularized MDPs with additional convexity in the primal variable, providing asymptotic convergence analysis using natural gradient methods, albeit under the assumption of full model knowledge. Neu et al. (2017) and Geist et al. (2019) formalize the entropy-regularized MDP framework, outlining the conditional entropy regularizer and the contraction properties of the regularized Bellman evaluation operator. Ying and Zhu (2020) offers a comprehensive overview of various LP formulations in both regularized and unregularized contexts, emphasizing their equivalence to Bellman equations and policy gradient methods. Furthermore, convex Q-learning (Meyn, 2022; Mehta and Meyn, 2020; Lu and Meyn, 2023) is motivated by the idea of deriving a convex relaxation of the squared Bellman error based on the LP approach. The method involves solving a constrained convex optimization problem over the primal variable, the Q-function. It is primarily designed for analyzing the non-tabular setting, where the value function is learned via function approximation, which distinguishes this approach from our own.
Finally, our finite-time analysis builds on recent last-iterate convergence-rate results for two-timescale stochastic approximation Zeng et al. (2024). We face additional challenges in our setting, namely asynchronous updates and replay-buffer gradient estimates that introduce bias. Despite this, on the high-probability good visitation event $\mathcal{G}_{\delta}$ (after a burn-in period), we prove a last-iterate mean-square rate $\widetilde{\mathcal{O}}(k^{-2/3})$ for the dual iterates, implying corresponding mean-square rates for the induced regularized value function and policy. Equivalently, achieving $\epsilon$ accuracy in mean-square error requires $k(\epsilon)=\widetilde{\mathcal{O}}(\epsilon^{-3/2})$ iterations to ensure $\mathbb{E}[\|\rho_{k}-\rho^{\star}\|_{2}^{2}\mid\mathcal{G}_{\delta}]\leq\epsilon$ (up to logarithmic factors).
@@
Structure. The structure of the paper is as follows. Section 2 formally introduces regularized MDPs and the associated saddle-point formulation that underlies our proposed algorithms. Section 3 presents a first algorithm in a synchronous setting, assuming access to a generative model. In Section 4, we extend this setting to the asynchronous case, where learning occurs from a single trajectory of data generated under a potentially changing policy. The section further contains the almost sure convergence result and our convergence rate analysis. Section 5 demonstrates the effectiveness of the proposed asynchronous algorithm on a standard reinforcement learning benchmark. Finally, Section 6 concludes the paper. The proofs details and technical results are deferred to the appendix to improve the readability of the main text.
@@
Notation. For any state $s\in\mathcal{S}$, we denote by $\delta_{s}$ the Dirac measure in $s$. For a variable $v(s,a)$ depending on state-action pairs $(s,a)\in\mathcal{S}\times\mathcal{A}$, we denote the state marginal by $\tilde{v}(s):=\sum_{a\in\mathcal{A}}v(s,a)$. For a finite set $\mathcal{S}$ with cardinality $n$, the probability simplex over $\mathcal{S}$ is denoted by $\Delta_{\mathcal{S}}=\{x\in\mathbb{R}_{+}^{n}:\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}x_{i}=1\}$. We use the symbol $e_{i}\in\mathbb{R}^{|\mathcal{S}|}$ for the $i$-th standard basis vector on $\mathbb{R}^{|\mathcal{S}|}$. For any logical expression $\mathcal{E}$, the indicator function $\mathbf{1}\{\mathcal{E}\}$ evaluates to $1$ if $\mathcal{E}$ is true and to 0 otherwise. To denote sequences, we use $\{\beta_{k}\}$ and $\{\beta(k)\}$ interchangeably for $k\in\mathbb{N}$. For $a,b\in\mathbb{R}$ we denote with $a\lor b$ the larger of $a$ and $b$ and with $\lfloor a\rfloor$ the integer part of $a$. For a compact set $H\subset\mathbb{R}^{d}$, $d\in\mathbb{N}$, we denote the Euclidean projection operator as $\Pi_{H}$, where $\Pi_{H}:\mathbb{R}^{d}\rightarrow H$ is defined by $\Pi_{H}(x)=\arg\min_{y\in H}\|x-y\|_{2}$. We denote the total variation norm as $\|\cdot\|_{\mathrm{TV}}$. For kernels $\mathcal{P},\,\mathcal{Q}$ on $\mathcal{X}$, we define $\|\mathcal{P}-\mathcal{Q}\|_{\infty}\;:=\;\sup_{x\in\mathcal{X}}\,\|\mathcal{P}(\cdot|x)-\mathcal{Q}(\cdot|x)\|_{\mathrm{TV}}.$
@@ 2 Problem Formulation
which admits a unique solution. We approach the unregularized MDP by its formulation as a linear program, Puterman (1994, Section 6.9). Reformulating the Bellman optimality equations as inequalities yields that a bounded function $V:\mathcal{S}\rightarrow\mathbb{R}$ that satisfies
@@
where $\mu\in\mathbb{R}_{+}^{|\mathcal{S}|}$ is any strictly positive vector such that $\sum_{s\in\mathcal{S}}\mu(s)=1$. It is well-known (Puterman, 1994, Section 6.9) that the optimizer to the primal LP (1) is exactly the optimal value function $V^{\star}_{ur}$. The corresponding dual linear problem is defined as
@@
where $\tilde{\rho}(s):=\sum_{a^{\prime}\in\mathcal{A}}\rho(s,a^{\prime})$ denotes the state marginal of the dual variable. Let $\rho^{\star}_{ur}$ denote a solution to the dual problem (2), then $\pi_{\rho_{ur}^{\star}}$ is an optimal policy, that is, $V^{\pi_{\rho^{\star}_{ur}}}_{ur}=V^{\star}_{ur}$, see (Puterman, 1994, Theorem 6.9.4). The Lagrangian formulation of the primal LP (1) is
@@ 2.1 Regularized MDP
Our approach is based on the regularized MDP formulation as formally described in Geist et al. (2019). Let $G:\mathbb{R}^{|\mathcal{A}|}_{+}\rightarrow\mathbb{R}$ be a strongly concave function and $\eta_{\rho}>0$ a regularization parameter. For a stationary policy $\pi\in\Pi$, the regularized Bellman operator $T_{\pi,G}:\mathbb{R}^{|\mathcal{S}|}\rightarrow\mathbb{R}^{|\mathcal{S}|},V\mapsto T_{\pi,G}V$ is defined state-wise as
@@
where the maximum is state-wise. Both operators are $\gamma$-contractions in the supremum norm, (Geist et al., 2019, Prop. 2). The regularized value function of a policy $\pi$, denoted $V_{r}^{\pi}$, is defined as the fixed point of the operator $T_{\pi,G}$. The regularized value function of the regularized MDP, denoted $V^{\star}_{r}$, is defined as the fixed point of the regularized optimality operator $T_{\star,G}$. Due to the strong concavity of $G$, the regularized Bellman optimality operator has a unique maximizing argument.
Proposition 1
(Geist et al., 2019, Theorem 1) The policy $\pi^{\star}_{r}:=\arg\max_{\pi\in\Pi}T_{\pi,G}V^{\star}_{r}$ is the unique optimal regularized policy in the sense that, state-wise, $V^{\pi^{\star}_{r}}_{r}=V^{\star}_{r}\geq V^{\pi}_{r}$ for all policies $\pi\in\Pi$.
The optimal regularized policy $\pi^{\star}_{r}$ is a strictly positive and exploratory softmax policy for entropy-type $G$. The Lagrangian formulation of the regularized MDP (Ying and Zhu, 2020) includes the strongly concave function of the dual variable $g(\rho):=\sum_{s\in\mathcal{S}}G(\rho_{s})$ with $\rho_{s}=\rho(s,\cdot)\in\mathbb{R}^{|\mathcal{A}|}_{+}$ and is given as
@@
The following corollary is a consequence of Proposition 1 and characterizes the solution to the regularized Lagrangian.
Corollary 2
@@
Proposition 3
(Geist et al., 2019, Theorem 2) Assume that constants $L_{G}$ and $U_{G}$ exist such that for all $\pi\in\Pi$ we have $L_{G}\leq\max_{s\in\mathcal{S}}G(\pi(\cdot|s))\leq U_{G}$. Then, $V^{\star}_{ur}(s)-\eta_{\rho}\frac{U_{G}-L_{G}}{1-\gamma}\leq V_{ur}^{\pi^{\star}_{r}}(s)\leq V^{\star}_{ur}(s)$ for all $s\in\mathcal{S}$.
For the Shannon entropy regularizer $G(\pi(\cdot|s)):=\sum_{a\in\mathcal{A}}\pi(a|s)\log(\pi(a|s)),$ we have $0\leq G(\pi(\cdot|s))\leq\log|\mathcal{A}|,$ for all $\pi\in\Pi$ and we can take $L_{G}=0,\ U_{G}=\log|\mathcal{A}|$ in the above stated bounds. To ensure better computational properties, we seek to make the regularized problem (7), defined by the regularized Lagrangian $L_{r}$, strongly convex in $V$. Since $L_{r}$ is originally linear in $V$, we propose an additional modification to problem (7). Specifically, we replace the linear term $\mu^{\top}V$ with a quadratic term and introduce a specific dual regularization function $g$ of the form
@@
The closely related recent work Li et al. (2024) also studies the convexified regularized MDP (9). They prove that the convex modification does not affect the optimal regularized value and only changes the optimal dual variable by a scaling factor that does not influence the optimal regularized policy.
Proposition 4
(Li et al., 2024, Theorem 2.1) The unique solution $(V^{\star},\rho^{\star})$ to the double-regularized problem (9) is such that $V^{\star}=V^{\star}_{r}$ and the optimal policy induced by $\rho^{\star}$ coincides with $\pi^{\star}_{r}$, that is, $\pi_{\rho^{\star}}=\pi_{\rho_{r}^{\star}}=\pi^{\star}_{r}$.
@@
Lemma 5 (Optimal dual variable)
The optimal dual variable of problem (9) with
bounded reward is strictly positive and bounded from above, that is, there exist $C^{L},C^{U}>0$ such that $C^{L}<\rho^{\star}(s,a)<C^{U}$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$.
Let $C^{L}$ and $C^{U}$ be the explicit bounds derived in the constructive proof of Lemma 5. Define $H$ as